# Zhang, Liu, and Zhao

### v3

@@ 10 Details of Convergence Analysis in Section 4
    Suppose that Assumptions 2 and 1 additionally hold with $M_{\Delta}>2\kappa q\mu^{2}$. It holds that $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}-\kappa]$.
Proof 10.3
Proof of Proposition 10.2. By Proposition 4.2, it suffices to prove the inclusion $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{*}$. Since the two max-affine components of $f(\cdot;\xi)$ for any $\xi\in S$ are Lipschitz continuous with modulus $L_{g,h}\coloneqq\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, we see that with $\varepsilon\geq\varepsilon_{0}=2L_{g,h}\sqrt{q}\mu$, the set $\mathcal{I}^{\varepsilon}(\theta^{\prime})$ contains all index mappings, that is, $\mathcal{I}^{\varepsilon}(\theta^{\prime})=\{(I_{1},I_{2})\mid I_{1}:S\rightarrow[K],I_{2}:S\rightarrow[K]\}$ for any $\theta^{\prime}\in\Theta$.
For the first statement, assume by contradiction that there exists $\widehat{\theta}^{d}\in\mathcal{S}^{d}_{\widehat{\varepsilon},\widehat{\rho}}\backslash\mathcal{S}^{*}$ for some $\widehat{\varepsilon}\in[\varepsilon_{0},+\infty)$ and $\widehat{\rho}\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$. Since the set $\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ contains all index mappings for the inner PA functions, by the surrogation construction for Case 1 and Case 2 in Appendix 9, for any $\theta^{*}\in\mathcal{S}^{*}$ there exists $I^{\widehat{\varepsilon}}\in\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ such that $\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})=\widehat{F}(\theta^{*};\widehat{\theta}^{d},I^{\widehat{\varepsilon}})=\min_{\theta\in\Theta}F(\theta)$. It follows that
@@
The following lemma from Deng and Gao (2021) is rephrased in our context for the proof of Proposition 4.3.
Lemma 10.4 (Deng and Gao (2021), Theorem 3.2)
Suppose that the stochastic function $\widehat{F}(\cdot;\theta^{\prime},\xi)$ is convex and Lipschitz continuous on $\Theta$ with modulus $L_{\widehat{F}}$ for any $\xi\in S$ and $\theta^{\prime}\in\Theta$, and that samples $\Xi_{N}\coloneqq\{\xi^{s}\}_{s=1}^{N}$ are i.i.d. generated from $S$. For $\eta>0$, denote $\theta^{\prime}_{+}\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{N}\sum_{s=1}^{N}\widehat{F}(\theta;\theta^{\prime},\xi^{s})+\frac{\eta}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}$. Then it holds that
$\left\lvert\mathbb{E}_{\Xi_{N}}\left[\frac{1}{N}\sum_{s=1}^{N}\widehat{F}(\theta^{\prime}_{+};\theta^{\prime},\xi^{s})-\mathbb{E}_{\widetilde{\xi}_{n}}\left[\widehat{F}(\theta^{\prime}_{+};\theta^{\prime},\widetilde{\xi}_{n})\right]\right]\right\rvert\leq\frac{2L_{\widehat{F}}^{2}}{\eta N}.$
Proof 10.5
Proof of Proposition 4.3. Let $\widetilde{I}^{\varepsilon}_{\nu}$ be a random mapping uniformly drawn from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, $\{\widetilde{\xi}^{s}_{\nu}\}_{s=1}^{N_{\nu}}$ be random vectors independently and uniformly drawn from $S$, $\mathcal{G}_{\nu-1}$ denote the $\sigma$-algebra generated by all past samples $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and mappings $\{\widetilde{I}^{\varepsilon}_{t}\}_{t=0}^{\nu-1}$, and $\mathcal{G}_{\nu-1}^{+}$ denote the $\sigma$-algebra generated by $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and $\{\widetilde{I}^{\varepsilon}_{t}\}_{t=0}^{\nu}$. Given the current iterate $\theta^{\nu}$ and the index mapping $I^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$, by the definition of $P_{I^{\varepsilon}_{\nu}}^{\rho}(\cdot)$ in (10), we have
$\min\left\{\widehat{F}(P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},I^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}+\frac{\rho}{2}\,\big{\lVert}\,{P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{\nu}).$ (17)
@@
$\frac{\rho}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$ $\leq F(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=\underbrace{F(\theta^{\nu})-W(\theta^{\nu})}_{\mathcal{E}_{1}^{\nu}}+\underbrace{W(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]}_{\mathcal{E}_{2}^{\nu}},$ (18)
where $W(\theta^{\nu})\coloneqq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$.
From the update rule in Step 5 of Algorithm 1, we have $\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\geq F_{N_{\nu}}(\theta^{\nu+1})$ and $F_{N_{\nu}}(\theta^{\nu})\geq F_{N_{\nu}}(\theta^{\nu+1})$. It follows that
$W(\theta^{\nu})$ $\geq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\mid\mathcal{G}_{\nu-1}\right]=\mathbb{E}\left[F_{N\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right].$
@@
$\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\{\phi(\cdot;\theta):S\rightarrow\mathbb{R}\mid\phi(\xi;\theta)=\varphi(f(\theta;\xi);\xi),\,\xi\in S,f(\theta;\cdot)\in{\mathcal{H}}_{\mu}^{K},\theta\in\Theta\}.$
From finiteness of $S$ and continuity of $\varphi(\cdot;\xi)$, it follows that for any $\phi\in\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}$, $\phi(\xi;\cdot)$ is Lipschitz continuous on the compact set $\Theta$ with uniform modulus, and that $\phi(\cdot;\theta)$ is uniformly bounded. Therefore, for the term $\mathcal{E}_{1}^{\nu}$, there exists a constant $C_{1}>0$ such that
$\mathcal{E}_{1}^{\nu}$ $\leq F(\theta^{\nu})-\mathbb{E}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
$=F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\sup_{\theta\in\Theta}\left\lvert F(\theta)-F_{N_{\nu}}(\theta)\right\rvert\,\middle|\,\mathcal{G}_{\nu-1}\right]$
@@
where the third inequality follows from inequalities (3.8)-(3.13) in Mohri et al. (2018) that are part of the proof of Lemma 8.5, and the last inequality is derived by Lemma 8.3.
Next, we bound the term $\mathcal{E}_{2}^{\nu}$. Given $I^{\varepsilon}_{\nu}$ and $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by the optimality of ${\theta}^{\nu+\frac{1}{2}}$, we have
$\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\leq\widehat{F}_{N_{\nu}}(P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}.$
@@
$W(\theta^{\nu})$ $\leq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right],$
where the equality holds since both $\theta^{\nu}$ and $P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})$ are independent of $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$. Then we obtain
$\mathcal{E}_{2}^{\nu}$ $\leq\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$ (20)
@@
$\frac{\rho-\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
@@
$\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\big{(}\,\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\theta^{\nu})]\,\big{)}^{2}\leq\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\mathbb{E}\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{0})-\mathbb{E}\left[F(\theta^{T})\right]+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
@@
$W(\theta^{\nu})=\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$

### v4

@@ 7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
    Suppose that Assumption (EC1) holds with $M_{\Delta}>2\kappa q\mu^{2}$, and that the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies Condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$. It holds that $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}-\kappa]$.
Proof 8.3
Proof of Proposition 8.2. By Proposition 3.4, it suffices to prove the inclusion $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{*}$. Since the two max-affine components of $f(\cdot,\xi)$ for any $\xi\in\Xi$ are Lipschitz continuous with modulus $L_{g,h}\coloneqq\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, we see that with $\varepsilon\geq\varepsilon_{0}=2L_{g,h}\sqrt{q}\mu$, the set $\mathcal{I}^{\varepsilon}(\bar{\theta})\coloneqq\{\{I^{\varepsilon}_{s}\}_{s=1}^{n}\mid I^{\varepsilon}_{s}\in\mathcal{I}(\bar{\theta},\xi^{s}),\,s=1,\dots,n\}$ contains all index combinations from the $\varepsilon$-active index sets $\{\mathcal{I}(\bar{\theta},\xi^{s})\}_{s=1}^{n}$, that is, $\mathcal{I}^{\varepsilon}(\bar{\theta})=[K]^{2n}$ for any $\bar{\theta}\in\Theta$.
For the first statement, assume by contradiction that there exists $\widehat{\theta}^{d}\in\mathcal{S}^{d}_{\widehat{\varepsilon},\widehat{\rho}}\backslash\mathcal{S}^{*}$ for some $\widehat{\varepsilon}\in[\varepsilon_{0},+\infty)$ and $\widehat{\rho}\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$. Since the set $\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ contains all index combinations for the inner PA functions, by the surrogation construction for Case 1 and Case 2 in Online Appendix 7, for any $\theta^{*}\in\mathcal{S}^{*}$ there exists $\{I^{\widehat{\varepsilon}}_{s}\}_{s=1}^{n}\in\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ such that $\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})=\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta^{*},\xi^{s};\widehat{\theta}^{d},I^{\widehat{\varepsilon}}_{s})=\min_{\theta\in\Theta}F(\theta)$. It follows that
@@
[…]
Suppose that a stochastic function $\phi(\cdot,\xi)$ is convex and Lipschitz continuous on $\Theta$ with modulus $L_{\phi}$ for any $\xi\in\Xi$ and $\bar{\theta}\in\Theta$, and that samples $\Xi_{N}\coloneqq\{\widetilde{\xi}^{s}\}_{s=1}^{N}$ are i.i.d. generated from $\Xi$. For $\eta>0$, denote $\bar{\theta}_{+}\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{N}\sum_{s=1}^{N}\phi(\theta,\widetilde{\xi}^{s})+\frac{\eta}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}$. Then it holds that
$\left\lvert\mathbb{E}_{\Xi_{N}}\left[\frac{1}{N}\sum_{s=1}^{N}\phi(\bar{\theta}_{+},\widetilde{\xi}^{s})-\mathbb{E}_{\widetilde{\xi}_{n}}\left[\phi(\bar{\theta}_{+},\widetilde{\xi}_{n})\right]\right]\right\rvert\leq\frac{2L_{\phi}^{2}}{\eta N}.$
Proof 8.5
Proof of Proposition 3.5. Let $\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}=\{\widetilde{I}^{\varepsilon}_{\xi}\}_{\xi\in\Xi}$ be a random index combination uniformly drawn from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, $\{\widetilde{\xi}^{s}_{\nu}\}_{s=1}^{N_{\nu}}$ be random vectors independently and uniformly drawn from $\Xi$. Let $\mathcal{G}_{\nu-1}$ denote the $\sigma$-algebra generated by all past samples $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and combinations $\{\widetilde{\mathbf{I}}^{\varepsilon}_{t}\}_{t=0}^{\nu-1}$, and $\mathcal{G}_{\nu-1}^{+}$ denote the $\sigma$-algebra generated by $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and $\{\widetilde{\mathbf{I}}^{\varepsilon}_{t}\}_{t=0}^{\nu}$. Given the current iterate $\theta^{\nu}$ and the index combination $\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$, by the definition of $P_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}^{\rho}(\cdot)$ in (15), we have
$\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}+\frac{\rho}{2}\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{\nu}).$ (25)
@@
$\frac{\rho}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$ $\leq F(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=\underbrace{F(\theta^{\nu})-W(\theta^{\nu})}_{\mathcal{E}_{1}^{\nu}}+\underbrace{W(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]}_{\mathcal{E}_{2}^{\nu}},$ (26)
where $W(\theta^{\nu})\coloneqq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{\nu}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$.
From the update rule in Step 5 of Algorithm 1, we have $\widehat{F}_{\nu}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\geq F_{\nu}(\theta^{\nu+1})$ and $F_{\nu}(\theta^{\nu})\geq F_{\nu}(\theta^{\nu+1})$. It follows that
$W(\theta^{\nu})$ $\geq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\mid\mathcal{G}_{\nu-1}\right]=\mathbb{E}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right].$
@@
$\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\{\phi(\cdot,\theta):\Xi\rightarrow\mathbb{R}\mid\phi(\xi,\theta)=\varphi(f(\theta,\xi),\xi),\,\xi\in\Xi,f(\theta,\cdot)\in{\mathcal{H}}_{\mu}^{K},\theta\in\Theta\}.$
From finiteness of $\Xi$ and continuity of $\varphi(\cdot,\xi)$, it follows that for any $\phi\in\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}$, $\phi(\xi,\cdot)$ is Lipschitz continuous on the compact set $\Theta$ with uniform modulus, and that $\phi(\cdot,\theta)$ is uniformly bounded. Therefore, for the term $\mathcal{E}_{1}^{\nu}$, there exists a constant $C_{1}>0$ such that
$\mathcal{E}_{1}^{\nu}$ $\leq F(\theta^{\nu})-\mathbb{E}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
$=F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\sup_{\theta\in\Theta}\left\lvert F(\theta)-F_{\nu}(\theta)\right\rvert\,\middle|\,\mathcal{G}_{\nu-1}\right]$
@@
[…]
Next, we bound the term $\mathcal{E}_{2}^{\nu}$. Given $\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}$ and $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by the optimality of ${\theta}^{\nu+\frac{1}{2}}$, we have
$\widehat{F}_{\nu}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\leq\widehat{F}_{\nu}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}.$
@@
$W(\theta^{\nu})$ $\leq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{\nu}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{\nu}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right],$
where the equality holds since both $\theta^{\nu}$ and $P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})$ are independent of $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$. Then we obtain
$\mathcal{E}_{2}^{\nu}$ $\leq\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$ (28)
@@
$\frac{\rho-\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
@@
$\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\big{(}\,\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\theta^{\nu})]\,\big{)}^{2}\leq\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\mathbb{E}\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{0})-\mathbb{E}\left[F(\theta^{T})\right]+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
@@
$W(\theta^{\nu})=\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$